can anyone help find the slope of the line containing the two given points? A. (5,9);(8,0) m= B. (1,5);(-1,-5) m= C. (0,-4);(3,-3) M= D. (6,4);(11,4) M= I just cant seem to be able to figure these out.
Answer by stanbon(75887)

find the slope of the line containing the two given points?
-------------------------------
General Rule: slope = [change in y] / [corresponding change in x]
-------------------------------
A. (5,9);(8,0) m= (0-9)/(8-5) = -9/3 = -3
---------------------------
B. (1,5);(-1,-5) m= (-5-5)/(-1-1) = -10/-2 = 5
---------------------------
C. (0,-4);(3,-3) M= (-3--4)/(3-0) = 1/3
---------------------------
D. (6,4);(11,4) M= (4-4)/(11-6) = 0/4 = 0
--------------------------
================================
Cheers,
Stan H.